# Boundary Attention: Learning to Find Faint Boundaries at Any Resolution

### 01 Jan 2024 v1

@@ Abstract
We present a differentiable model that explicitly models boundaries—including contours, corners and junctions—using a new mechanism that we call boundary attention. We show that our model provides accurate results even when the boundary signal is very weak or is swamped by noise. Compared to previous classical methods for finding faint boundaries, our model has the advantages of being differentiable; being scalable to larger images; and automatically adapting to an appropriate level of geometric detail in each part of an image. Compared to previous deep methods for finding boundaries via end-to-end training, it has the advantages of providing sub-pixel precision, being more resilient to noise, and being able to process any image at its native resolution and aspect ratio.
@@ 1 Introduction
Converting a precise contour that is defined mathematically in continuous 2D space to its discrete representation in pixel-space is a common task in computer graphics, and there are established techniques for rasterization [12, 27], anti-aliasing [16, 6] and so on. However, the inverse problem in computer vision of robustly inferring precise and unrasterized descriptions of contours from discrete images remains an open challenge. Earlier work in the field inferred unrasterized representations of edges [4, 8, 25, 14, 13, 22]—and sometimes corners and junctions too—and it explored algorithms for connecting them into boundaries via edge-linking and MRFs. But since the dawn of deep learning, boundaries have almost exclusively been modeled using discrete, rasterized maps. Moreover, it is not uncommon to rely heavily on supervision from human annotations, and even modern datasets can suffer from annotation errors on the order of pixels, quite far from sub-pixel accuracy.
Taking inspiration from early computer vision work, we present a model for inferring unrasterized boundaries that can be learned instead of designed. We show that this model is able to benefit from the power of deep learning while also realizing many of the advantages—such as robustness to noise, sub-pixel precision, and adaptability between signal-types—that have long been the potential strengths of classical bottom-up techniques.
@@
The core of our model is a mechanism we call boundary attention. It is a boundary-aware local attention operation that, when applied densely and repeatedly, progressively refines a field of variables that specifies the local boundaries surrounding every pixel. The model’s output is a field of overlapping geometric primitives that can be used in a variety of ways, including to produce an unsigned distance function for the image’s boundaries, a boundary-aware smoothing of its channel-values, and a field of soft local attention maps that relate every pixel to its neighbors.
To evaluate our model, we consider the problem of finding boundaries in images corrupted by extreme amounts of noise [23, 24, 33], where the images can be of any size and resolution, and where we know little about the noise and the objects or shapes we are looking for. We choose this problem because it relies entirely on having a strong model for the basic topological and geometric properties of boundaries, namely that they are piecewise smooth curves which connect at corners or junctions and divide regions of homogeneity.
Figure 1: Our model uses neighborhood attention with dense, stride-1 tokens (but we depict non-overlapping ones here for visual clarity). The entire model is invariant to discrete spatial shifts and so applies at any resolution. Each token encodes an adaptively-sized geometric primitive that represents the unrasterized local boundaries surrounding a pixel, and the tokens evolve through boundary attention to be geometrically consistent. The output is an overlapping field of primitives that directly implies a boundary-aware smoothing of the input image and an unsigned distance map to the image boundaries.
@@
We show that our model has several advantages. All of its components are local and invariant to discrete spatial shifts, so it can be trained on small-sized images and then deployed on much larger and differently-shaped ones. It is also very compact, comprising only 207k parameters, and it runs several times faster than many alternatives. We also find that it can be trained to a useful state with very simple synthetic data, made up of random circles and triangles that are uniformly colored and then corrupted by noise. Despite this simplicity, we find that the model can generalize to real images surprisingly well.
@@
    We demonstrate that our model based on boundary attention is more effective than current state-of-the-art methods at finding boundaries at high levels of sensor noise (Fig. LABEL:fig:real_results) and can achieve sub-pixel accuracy, even in the presence of noise.
@@ 2 Related Work
It is possible to process the outputs of certain filters to localize edges with sub-pixel precision (e.g.,  [4, 8, 25, 14, 13, 22]). But this approach struggles near corners and junctions because the edge filters are not derived for these, and accuracy breaks down. Recently, the field of junctions [33] showed improvements by expanding the library of local, unrasterized geometric primitives from classical edge-based primitives to a larger family of “generalized junctions”. This allows unifying a variety of bottom-up cues that had often previously been handled separately, such as geometrically-consistent and polarity-consistent edge elements being linked into contours [18, 23, 24, 4]; contours that approach one another being joined at corners and junctions [15, 21, 36]; homogeneous regions that look the same being together; and those that look different being apart [7, 11, 2]. The field of junctions showed that putting all of these cues together leads to a dramatic improved in noise resilience, and it also allows analyzing the recovered boundaries explicitly into its component junctions, edges and corners. In this paper, we leverage all of these benefits, and we do this more efficiently and with an adaptive patch size. Additionally, our model is differentiable, meaning it can be used as a component in larger systems where accurate boundaries are necessary.
Our work is complementary to recent approaches that leverage large-scale training to perform edge detection or segmentation by internalizing the patterns in a training set. Recent examples include Segment Anything [19] and combinations of strong self-supervised features with clustering and CRFs [5, 3, 20]. These large-scale approaches rely on large training sets and high-capacity architectures that can internalize high-level cues like object and scene familiarity as well as low-level cues like spatial coherence. Our approach is different because it relies exclusively on low-level cues, has much less capacity, and uses much less training data. Despite this difference, we find that our model’s inherent preference for contour-like boundaries allows it to achieve comparable (or better) results, in particular when noise is high. This suggests that future work could benefit from incorporating some of our architectural elements into larger and higher-capacity systems, in particular to improve interpretability, efficiency, generalization or task-transfer.
@@ 3 Representation
Our system is depicted in Figure 1. It uses neighborhood cross-attention, a patch-wise variant of cross-attention, with pixel-resolution tokens. Critically, each pixel’s token is explicitly required to encode a tuple of values representing a geometric primitive that specifies the geometric structure of the local boundaries surrounding the pixel. Because the tokens are forced through this geometric bottleneck, we refer to our model’s core mechanism as Boundary Attention.
We instantiate our bottleneck using a learned linear mapping from the token dimension to a predefined lower-dimensional space of unrasterized boundary patterns that we call junction space. Our choice for this space is inspired by the Field of Junctions [33], but we use a modified parameterization (described below) that is differentiable across its entire domain. Junction space has the benefit of specifying local boundary patterns without rasterization and thus with unlimited spatial precision. As depicted in Figure 2 and described in [33], it also has the benefit of including a large family of local boundary patterns, including uniformity (i.e., absence of boundaries), edges, bars, corners, T-junctions and Y-junctions.
The next section introduces our parameterization of junction space and some associated operators. Then Section 4 describes the architecture that we use to analyze an image into its field of junction values.
Figure 2: Samples on a smooth trajectory through our space of geometric primitives (“junction space”), with one sample enlarged and accompanied by a visualization of its distance map on the right. Each sample comprises a vertex position $(u,v)$, orientation $\theta$, and angles $(\omega_{1},\omega_{2},\omega_{3})$ that partition the circle. The space includes all types of locally-uniform regions, edges, bars, corners, Y-junctions and T-junctions.
@@ 3.1 Boundary Primitives
Denote the $K$-channel input image by $\{\mathbf{f}[n]\}$, where $\mathbf{f}[n]\in\mathbb{R}^{K}$ is the vector image value at the discrete pixel grid index $n$. Our approach is to analyze the image into a field of dense, stride-1 overlapping local patches, each having a square support $\Omega_{n}(x)$ centered at the $n$th pixel.
@@
There are many ways to partition a local region $\Omega_{n}(x)$, and one can define parametric families of such partitions. For example the set of oriented lines provides a two-parameter family of partitions, with each member of the family separating the region into points that lie on one side of a line or the other. This family of partitions would be appropriate for describing edges. Here we define a larger family of partitions that encompasses a greater variety of local boundary structures.
@@
As depicted in the right of Figure 2, our partitions are parameterized by $\mathbf{g}\in\mathbb{R}^{2}\times\mathbb{S}^{1}\times\triangle^{2}$, where $\mathbb{S}^{1}$ is the unit circle and $\triangle^{2}$ is the standard $2$-simplex. We use the notation $\mathbf{g}=(\boldsymbol{u},\theta,\boldsymbol{\omega})$, where $\boldsymbol{u}=(u,v)\in\mathbb{R}^{2}$ is the vertex, $\theta\in\mathbb{S}^{1}$ is the orientation, and $\boldsymbol{\omega}=(\omega_{1},\omega_{2},\omega_{3})$ are barycentric coordinates (defined up to scale) for the three relative angles, ordered clockwise starting from $\theta$. Our convention is to express the vertex coordinates relative to the center of region $\Omega_{n}(x)$, and we denote that the vertex is free to move outside of this region. We also note that up to two angles $\omega_{j}$ can be zero. This all makes it possible to represent a variety of partition types, including edges, bars, corners, 3-junctions and uniformity (i.e., trivial or singleton partitions).
@@
Fixing a value for $\mathbf{g}$ induces three binary-valued spatial support functions:
$s_{nj}(x;\mathbf{g})\in\{0,1\},\ j=1,2,3,$ (1)
@@
that evaluate to $1$ for points $x$ that are in $\Omega_{n}(x)$ and in the $j$th wedge defined by $\mathbf{g}$; and 00 otherwise. It also induces an unsigned distance function:
$d_{n}(x;\mathbf{g})\geq 0,$ (2)
that represents the Euclidean distance from point $x$ to the nearest point in the boundary set defined by $\mathbf{g}$.
Defining these as continuous functions allows specifying the local boundary structure with unlimited spatial resolution. The right of Figure 2 uses three colors to visualise the wedge supports of a junction $\mathbf{g}$, and it shows a quantized visualization of the associated distance function.
Instead of fixing the size of every patch in the image to a hand-selected value that is the same throughout, we would like the size of every region $\Omega_{n}$ to adapt to the local geometry. We do this by equipping each patch with a parameterized local windowing function $w_{n}(x;\mathbf{p})\in[0,1]$. Specifically, we parameterize them as convex combinations of $W$ square window function coefficients $\mathbf{p}\in{\cal P}=\triangle^{W-1}$. That is,
$w_{n}(x;\mathbf{p})=\sum_{i=1}^{W}p_{i}\mathbf{1}[\|x-c[n]\|_{\infty}\leq D_{i}],$ (3)
where $\|\cdot\|_{\infty}$ is the $\ell^{\infty}$-norm, and $\mathbf{1}[\cdot]$ is the indicator function that returns $1$ if the argument true; and 00 otherwise. In our experiments we use $W=3$ and diameters $\mathbf{D}=(3,9,17)$.
Figure 3: Visualizing our model’s output. Top row, left to right: (i) input image; (ii) boundary-aware smoothing; and (iii) boundary map with attention maps at four query points. The two bottom rows show regions $\Omega_{n}$ unfolded from within the two white windows indicated above. Bottom rows, left to right: (i) window functions; (ii) boundaries; and (iii) support functions colored by their gathered wedge features. See text for details.
@@ 3.2 Gather and Slice Operators
Our network operates by refining the field $\{(\mathbf{g}^{t}[n],\mathbf{p}^{t}[n])\}$ over a fixed sequence of steps $t=1,...,T$. It uses two operators that we define here and depict in the right of Figure 4. The first operator is a patch-wise gather operator, in which each wedge of each region computes the weighted average of the image values it contains (recall that $c[n]$ are $n$th pixel’s coordinates):
$\mathbf{f}_{kj}=\frac{\sum_{n}\mathbf{f}[n]w_{k}(c[n];\mathbf{p}[n])s_{kj}(c[n];\mathbf{g}[n])}{\sum_{n}w_{k}(c[n];\mathbf{p}[n])s_{kj}(c[n];\mathbf{g}[n])}.$ (4)
The second operation is a pixel-wise slice operation, where each pixel computes the means and variances, over all regions that contain it, of the per-region distance maps $d_{n}(x;\mathbf{g}[n])$ and gathered wedge features $\mathbf{f}_{kj}$. The expressions for the means are:
@@
Note that the only contributions to the sums over patches $k$ are from patches that contain $c[n]$, i.e., the sums are over $\{k\mid\Omega_{k}\ni c[n]\}$. The expressions for the pixel-wise distance map variance $\nu_{d}[n]$, and for the feature variance $\nu_{f}[n]$, which is computed across patches containing $n$ and across their $K$ channels, are defined accordingly and included in the supplement.

### 18 Mar 2024 v2

# Boundary Attention: Learning to Localize Boundaries under High Noise

@@ Abstract
We present a differentiable model that infers explicit boundaries, including curves, corners and junctions, using a mechanism that we call boundary attention. Boundary attention is a boundary-aware local attention operation that, when applied densely and repeatedly, progressively refines a field of variables that specify an unrasterized description of the local boundary structure in every overlapping patch within an image. It operates in a bottom-up fashion, similar to classical methods for sub-pixel edge localization and edge-linking, but with a higher-dimensional description of local boundary structure, a notion of spatial consistency that is learned instead of designed, and a sequence of operations that is end-to-end differentiable. We train our model using simple synthetic data and then evaluate it using photographs that were captured under low-light conditions with variable amounts of noise. We find that our method generalizes to natural images corrupted by real sensor noise, and predicts consistent boundaries under increasingly noisy conditions where other state-of-the-art methods fail.
@@ 1 Introduction
Converting a precise contour that is defined mathematically in continuous 2D space to a discrete pixel representation is a common task in computer graphics, and there are established tools for rasterization [8, 18], anti-aliasing [11, 4] and so on. However, the inverse problem in computer vision of robustly inferring precise, unrasterized contours from discrete images remains an open challenge, especially in the presence of noise.
Earlier work by Canny [3] and many others [5, 16, 10, 9, 15] explored the detection of parametric edge models, and there are a variety of bottom-up algorithms that try to connect them by encouraging geometric consistency via edge-linking or message-passing. But since the dawn of deep learning, boundaries have almost exclusively been represented using discrete, rasterized maps, and spatial-consistency mechanisms that were previously based on explicit curve geometry have largely been replaced by black-box neural networks.
In this paper, we take inspiration from early computer vision work and revisit the task of finding unrasterized boundaries via bottom-up geometric processing. But unlike early work, we want to leverage the speed and power of neural networks to learn these processes instead of hand-crafting them.
We focus our attention on the low-level task of finding boundaries that separate regions of uniform color, as depicted by the toy problem in Figure 1. This task becomes difficult at high noise levels, especially within local patches, and we address it by creating a model that can learn to exploit a wide range of low-level cues, such as curves being predominantly smooth with low-curvature; curves tending to meet at corner and junction points; contours tending to have consistent contrast polarity throughout their extent; colors tending to vary smoothly at locations away from boundaries; and so on.
@@
The core of our model is a mechanism we call boundary attention. It is a boundary-aware local attention operation that, when applied densely and repeatedly, progressively refines a field of variables that specify the local boundaries within dense (stride-1) patches of an image. The model’s output is a dense field of unrasterized geometric primitives that, as depicted in the right of Figure 1, can be used in a variety of ways, including to produce an unsigned distance function and binarized map for the image boundaries, a boundary-aware smoothing of the input image colors, and a map of spatial affinities between any query point and the pixels that surround it.
An important feature of our model is that its output patch primitives (see Figure 2) have enough flexibility to represent a wide range of local boundary patterns and scales, including thin bars, edges, corners and junctions, all with various sizes, and all without rasterization and so with unlimited resolution. This gives our model enough capacity to learn to localize boundaries with high precision, and to regularize them in the presence of noise without erroneously rounding their corners or missing their junction points.
Figure 1: Process overview. The image unfolds into stride-1 patches, and boundary attention operates iteratively on their embeddings to produce for each patch: (i) a parametric three-way partitioning, and (ii) a parametric windowing function that defines its effective patch size. (Figure 2 shows parameterization details.) This output field implies a variety of global maps, shown in clockwise order: a boundary-aware smoothing of the input colors; an unsigned boundary-distance map; a boundary map; and a map of spatial affinities between any query point and its neighbors.
@@
We design our model so that all of its components are local and invariant to discrete spatial shifts, enabling it to be trained on small-sized images and then deployed on much larger and differently-shaped ones. By tying the weights across multiple portions of the network, we increase its resilience to noise which we validate with ablations. Our resulting model is very compact, comprising only 207k parameters, and it runs several times faster than many alternatives. Further, our model can be trained to a useful state with very simple synthetic data, made up of random circles and triangles that are uniformly colored and then corrupted by noise. Despite this simplicity, we find that our model can generalize to captured photographs and novel noise processes surprisingly well. We evaluate our model on real noisy images and show that it can capture fine details that other methods miss, and moreover maintain its predictions over high levels of noise.
@@
    We demonstrate that our model based on boundary attention is more effective than current state-of-the-art methods at precisely and consistently finding boundaries at high levels of sensor noise (Figure 6).
Figure 2: Parameterization details. Left: Each patch $k$ is associated with an unrasterized three-way partitioning of its area (colored blue, orange and purple here). The partitioning parameters comprise a vertex $(u,v)$, orientation $\theta$, and angles $(\omega_{1},\omega_{2},\omega_{3})$, defined up to scale. A: A walk through junction space by linearly interpolating between junctions is spatially smooth, and can represent edges, bars, corners, Y-junctions, T-junctions and uniform regions. B: Each junction is modulated through a learned windowing function. The windowing parameters $\mathbf{p}=(p_{1},p_{2},p_{3})$ are convex weights over a dictionary of binary pillboxes. See the supplement for an interactive visualization.
@@ 2 Related Work
Early approaches to edge detection rely primarily on low-level or local information. Canny [3] and others (e.g.[5, 16, 10, 9]) use carefully chosen filters, and later ones like pB [15] and gPb [14] combine such filters using a handful of trainable parameters. These methods provide local estimates of straight-edge position and orientation and they can subsequently be filtered and joined using geometry-based processes such as edge-linking [3]. However, they often struggle near corners and junctions due to the difficulty of designing filters for these more complicated structures. Structured edges [7] and other early learning-based methods [6, 13] are able to detect more-complicated structures, but they represent their local outputs using discrete, rasterized boundary maps that are not directly compatible with continuous boundary parameterizations or geometry-based linking.
Recently, the field of junctions (FoJ) [25] introduced a way to detect more-complicated local structures while also representing them parametrically, without rasterization. By using its unrasterized representations in a customized non-convex optimization routine, the method can detect curves, corners and junctions with high precision and with unprecedented resilience to noise. Our model is strongly inspired by this work, and it improves upon it by introducing an enhanced family of parameterized geometric primitives (Figure 2) and by replacing its hand-designed objective and associated non-convex optimization routine with a learned model that is fast, feed-forward and differentiable. We also tackle the challenge of representing images that vary in scale across different areas. Unlike many low-level methods that assess image edges at a single, frequently manually chosen scale, our approach considers variations in scale throughout the image. The field of junctions, for example, requires setting a single patch size that junctions are optimized over, where larger patch sizes allow for better resilience to noise at the expense of memory and detail preservation. In contrast we equip each local region with a function to modulate its spatial extent.
Our work is very different from the recent trend in computer vision of using deep networks to detect boundaries that are semantically meaningful. Examples include HED [27], EDTER [19], and UAED [28].
These models are trained using annotated segmentation datasets like BSDS500 [2] and NYUDv2 [22], and they aim to identify boundaries between semantic objects (e.g., people, animals) while ignoring boundaries that are inside of these objects (e.g., stripes on a shirt). Their large capacities and their end-to-end training allow them to exploit not only low-level cues but also high-level cues such as object and scene familiarity.
In contrast to this recent trend, we follow earlier work by focusing entirely on identifying boundaries from the bottom up, using low-level cues alone. Since our model is trained using simple synthetic data, it does not exploit object familiarity or learn to ignore intra-object edges. We find that this approach maintains some advantages: instead of producing a pixelized boundary map, our model produces a rich field of unrasterized primitives that provide better precision as well as information about local grouping. Moreover, despite not emphasizing top-down information, we find it performs better than deep models in the presence of high noise levels. This suggests that future work may benefit from incorporating some of our architectural elements into larger and higher-capacity systems, in particular to improve interpretability, efficiency, generalization and task-transfer.
@@ 3 Representation
Our system is depicted in Figure 1. It uses neighborhood cross-attention, a patch-wise variant of cross-attention, with stride-1 patch embeddings. Critically, each embedding explicitly encodes a tuple of values that specifies the geometric structure and spatial extent of the local boundaries within a patch. Because the embeddings are forced through this geometric bottleneck, we refer to our model’s core mechanism as boundary attention.
We instantiate our bottleneck using a learned linear mapping from the embedding dimension to a predefined lower-dimensional space of unrasterized boundary patterns that we call junction space. Junction space has the benefit of specifying local boundary patterns without rasterization and thus with unlimited spatial precision. As depicted in Figure 2 and described in [25], it also has the benefit of including a large family of local boundary patterns, including uniformity (i.e., absence of boundaries), edges, bars, corners, T-junctions and Y-junctions. We also modulate the spatial extent of each junction by learning an associated windowing function, described in Section 3.1.
To enable communication between patch embeddings, each of which corresponds to a local region of an image, we leverage the idea that each region of the image is described by multiple patches, and that these patches should agree on compatible representations of the global image structure. The ideal embedding for each patch should aptly represent its local geometry while agreeing with the structural claims of neighboring patches.
The next section introduces our parameterization of junction space and some associated operators. Then Section 4 describes the architecture that we use to transform an image into its junction space representation.
@@ 3.1 Boundary Primitives
Denote the $K$-channel input image by $\{\mathbf{f}[n]\}$, where $\mathbf{f}[n]\in\mathbb{R}^{K}$ is the vector image value at the discrete pixel grid index $n$. Our approach is to analyze the image into a field of dense, stride-1 overlapping local patches, each having a square support $\Omega_{k}(x)$ for the $k$th patch centered at the $k$th pixel.
@@
There are many ways to partition a local region $\Omega_{k}(x)$, and one can define parametric families of partitions. For example the set of oriented lines provides a two-parameter family of partitions, with each member of the family separating the region into points that lie on one side of a line or the other. This family of partitions would be appropriate for describing edges. Here we define a larger family of partitions that encompasses a greater variety of local boundary structures.
@@
As depicted in the right of Figure 2, our partitions are parameterized by $\mathbf{g}\in\mathbb{R}^{2}\times\mathbb{S}^{1}\times\triangle^{2}$, where $\mathbb{S}^{1}$ is the unit circle and $\triangle^{2}$ is the standard $2$-simplex. We use the notation $\mathbf{g}=(\boldsymbol{u},\theta,\boldsymbol{\omega})$, where $\boldsymbol{u}=(u,v)\in\mathbb{R}^{2}$ is the vertex, $\theta\in\mathbb{S}^{1}$ is the orientation, and $\boldsymbol{\omega}=(\omega_{1},\omega_{2},\omega_{3})$ are barycentric coordinates (defined up to scale) for the three relative angles, ordered clockwise starting from $\theta$. Our convention is to express the vertex coordinates relative to the center of region $\Omega_{k}(x)$, which is located at $c[k]$, and we denote that the vertex is free to move outside of this region. We also note that up to two angles $\omega_{j}$ can be zero. This all makes it possible to represent a variety of partition types, including edges, bars, corners, 3-junctions and uniformity (i.e., trivial or singleton partitions), and do so with unlimited spatial resolution.
@@
Fixing a value for $\mathbf{g}$ induces three wedge support functions, denoted by
$s_{kj}(x;\mathbf{g})\in\{0,1\},\ j=1,2,3.$ (1)
@@
These evaluate to $1$ for points $x$ that are in $\Omega_{k}(x)$ and in the $j$th wedge defined by $\mathbf{g}$; and 00 otherwise. It also induces an unsigned distance function, denoted by
$d_{k}(x;\mathbf{g})\geq 0,$ (2)
which represents the Euclidean distance from point $x$ to the nearest point in the boundary set defined by $\mathbf{g}$. Figure 2 uses three colors to visualise the wedge supports $s_{kj}$ of a junction $\mathbf{g}$, and Figure 3 shows the unsigned distance functions for $3\times 3$ grids of junction parameters. Analytic expressions for them are included in the supplement.
In order to enable the size of each region $\Omega_{n}$ to be able to adapt to the local geometry and noise conditions, we equip each one with a parameterized local windowing function $w_{k}(x;\mathbf{p})\in[0,1]$, with parameters $\mathbf{p}\in{\cal P}=\triangle^{W-1}$ that are the coefficients of a convex combination of $W$ square pillbox functions. That is,
$w_{k}(x;\mathbf{p})=\sum_{i=1}^{W}p_{i}\mathbf{1}[\|x-c[k]\|_{\infty}\leq D_{i}],$ (3)
[…]
Figure 3: Example of our model’s output, with examples from two different regions. Top row: Some of each region’s overlapping input patches, their corresponding outputs (visualized in the style of Figure 2), and three types of per-patch attributes that the outputs imply: unsigned distance; boundaries; and gathered wedge colors. Bottom row: Four types of global maps that are implied by accumulating values from the output field and rendered patches. See the supplement for interactive visualizations.
@@ 3.2 Gather and Slice Operators
Our network operates by refining the field $\{(\mathbf{g}^{t}[k],\mathbf{p}^{t}[k])\}$ over a fixed sequence of steps $t=1,...,T$. It uses two operators that we define here and depict in the right of Figure 4 to facilitate intra-patch communication between pixels within a single patch, and inter-patch communication across different patches representing overlapping image regions. The first operator is a patch-wise gather operator, in which each wedge of each patch computes the weighted average of the image values it contains (recall that $c[n]$ are the $n$th pixel’s coordinates, and wedges are indexed by $j$):
$\mathbf{f}_{kj}=\frac{\sum_{n}\mathbf{f}[n]w_{k}(c[n];\mathbf{p}[k])s_{kj}(c[n];\mathbf{g}[k])}{\sum_{n}w_{k}(c[n];\mathbf{p}[k])s_{kj}(c[n];\mathbf{g}[k])}.$ (4)
The second operation is a global pixel-wise slice operation, where each pixel computes the means and variances, over all regions that contain it, of the per-region distance maps $d_{k}(x;\mathbf{g}[k])$ and gathered wedge features $\mathbf{f}_{kj}$. The expressions for the means are:
@@
Note that only patches that contain $c[n]$ contribute to the sum, i.e., the sums are over $\{k\mid\Omega_{k}\ni c[n]\}$. The expressions for the pixel-wise distance map variance $\nu_{d}[n]$, and for the feature variance $\nu_{f}[n]$, which is computed across patches containing $n$ and across their $K$ channels, are defined accordingly and included in the supplement. Intuitively, slicing $\mathbf{f}_{kj}$ can be thought of globally accumulating information from each patch’s junction representation.